OpenStudy (anonymous):

The wavelength of red light is given as 6.5 × 10-7 m. If another wave of an unknown type has a frequency of 5.3 × 1015 Hz. What is the relation between the wavelength of the unknown wave and red light? (note: c = 2.998 × 108 m/s)

OpenStudy (shikamaru11):

by using relation v=fx lambda the wavelength of the unknown wave comes out to be 5.88 x 10-8m hope it helped @Mrs.SquarePants

OpenStudy (michele_laino):

the wavelength of the unknown radiation is: \[\lambda = \frac{c}{\nu } = \frac{{3 \cdot {{10}^{10}}}}{{5.3 \cdot {{10}^{15}}}} = ...{\text{cm}}\]